# From Conversation to Automation: Leveraging Large Language Models to Analyze Strategies in Problem Solving Therapy

### v1

@@ 6 Conclusion
This study highlights the potential of LLMs like GPT-4o to enhance PST through effective annotation of therapeutic dialogues. GPT-4o achieved a notable F1 score of 0.76 in the ‘No Context’ condition, with a low entropy value of 0.035, demonstrating its efficiency and reliability in categorizing therapeutic strategies. Additionally, we fine-tuned transformer-based models on a sample of 5,000 therapist utterances annotated by GPT-4o. Notably, DeBERTa matched the performance of our second-best LLM, GPT-4, showcasing the effectiveness of integrating AI with human expertise. We also expanded our codebook to include a new ‘Communication’ dimension, enriching our analysis of therapist-client interactions. Our findings showed that all models performed better at identifying ‘PS core’ strategies, indicating a strong capability in recognizing structured therapeutic elements. This research sets the stage for future innovations that enhance the integration of technology in mental health care, making therapeutic interventions more accessible and impactful.
@@ 7 Limitations
Our study has several limitations, including the dependence on ‘transcribed’ therapy sessions, which may not fully capture the dynamic aspects of live interactions such as tone and non-verbal cues. The use of LLMs like GPT-4o and GPT-4 introduces potential biases from their internet-based training data, impacting the interpretation of therapeutic strategies. Furthermore, our analysis is based solely on English-language data, which limits the model’s applicability in non-English-speaking contexts. These limitations highlight the technology’s current stage and its focus on textual analysis, excluding the broader emotional and psychological facets of therapy sessions.
@@
Finally, we did not include any analysis of LLM-generated therapy data, which could offer a valuable comparison between real-world therapeutic interactions and simulated data produced by LLMs. Our focus is exclusively on the characteristics of authentic, real-world data from therapy sessions. Future research could explore this comparison, examining how well LLM-generated interactions align with or diverge from real-world therapeutic PST dialogues, potentially offering insights into improving the practicality and effectiveness of AI-driven interventions in mental health care.
@@ 8 Ethical Considerations
A few of the coauthors provided the deidentified transcriptions of PST sessions, which had been collected as part of a large trial assessing the comparisons across different modalities of PST delivery. We used these transcriptions to conduct secondary analysis and experiment on automated annotation, as we have presented in this paper. The institutional review boards have determined this study as not human subjects research, and a data-sharing agreement has been signed.
The application of LLMs in healthcare, particularly in sensitive areas like cognitive behavioral therapy, necessitates careful consideration of their implications. Ensuring responsible development and use of AI technologies in mental health care is crucial, requiring vigilance for biases, stakeholder engagement, and strong safeguards to maintain individuals’ confidentiality and the integrity of the therapeutic process. This paper does not endorse the use of LLMs in therapeutic settings, nor does it claim that they are ready for such applications. Rather, our goal is to systematically evaluate the behavior of current LLMs when applied in therapeutic contexts. The use of AI in mental health care raises ethical concerns about automating therapy. AI tools, like LLMs, should assist rather than replace therapists, ensuring they do not undermine the personal, empathetic nature of therapy, which is crucial for effective treatment. While AI can provide useful insights, therapists must remain the primary decision-makers, as AI cannot fully grasp the complexities of human behavior and emotions as shown in our results and other related work. Accountability should always rest with human professionals to preserve the integrity and effectiveness of mental health care.
@@ Appendix B Therapists Strategies
Problem-solving Positive Mindset Strategies building a positive approach towards challenges, emphasizing optimism and emotional awareness. “uh, they suggest trying to think of problems as challenges rather than threats.”
Defining Problems and Goals Setting achievable goals and identifying factual problems while recognizing obstacles. “Are there any obstacles to those goals? Can we list anything that is getting in the way?”
PS Core Generating Alternative Solutions Brainstorming solutions, differentiating between strategies, and encouraging open idea exploration. “So let’s brainstorm some specific strategies and tactics for how you can make that happen.”
Evaluating Outcomes and Developing a Plan Narrowing down solutions, assessing impacts, and creating an action plan. “Working through this worksheet…short-term personal consequences of self-care are pretty positive, right?”
Trying Out Solution Plan Implementing and assessing the solution strategy, troubleshooting, and acknowledging successes. “Then just monitoring your outcomes. Has the problem improved in some ways?”
@@
Communication Session Management* The therapist provides guidance on session logistics, manages time, keeps discussions on track, and helps plan future sessions. “So you got Skype up and running, and that is so fun!”
@@
Problem-Solving Test Review* The therapist explains the PS Test and reviews the client’s results across five dimensions: positive/negative problem orientation, rational, impulsive, and avoidant problem-solving styles. “So positive and rational were your highest scores. Your positive was, uh, really, is almost as high as it could possibly get.”
@@ Appendix C Model Setting
Adam optimizer, learning-rate=2e-5, DeBERTa = microsoft/deberta-v3-base, MentalBERT = mental/mental-bert-base-uncased, FlanT5 = google/flan-t5-base
@@ Appendix D Prompts
D.1 Without Context
@@
Your Task: Select one strategy from the ‘PS core’ category and one strategy from the ‘Communication’ category that best aligns with the therapist’s utterance. If none of the strategies are applicable, return ’None’ for that category.
@@
Communication Strategies: 6- Social Courtesies: Engaging in polite, yet somewhat superficial social interactions like greetings, small talk (e.g., asking about the weather or weekend plans), thanking the client for their time, and saying goodbye. 7- Session Management: Managing session logistics, such as informing the client about session length, guiding the direction of the conversation, refocusing when off-topic, or scheduling future sessions. 8- Therapeutic Engagement: Actively listening, validating the client’s feelings, providing emotional support, recalling previous information shared by the client, reflecting emotions, normalizing experiences, commenting on client strengths, offering personal disclosure, and offering comfort. 9- Problem-solving Review Test: Discussing the client’s Problem-Solving Test (PST) results, explaining the test, and reviewing five problem-solving dimensions: positive problem orientation, negative problem orientation, rational style, impulsive style, and avoidant style.
D.2 With Context
@@
Your Task: Select one strategy from the ‘PS core’ category and one strategy from the ‘Communication’ category that best aligns with the therapist’s most recent utterance. You are also provided with the last two utterances: the therapist’s previous utterance and the client’s previous utterance. You should use only these two previous utterances as context, but for annotation, you should focus only on the most recent therapist’s utterance. If none of the strategies are applicable, return ’None’ for that category.
@@ E.1 LLM Classification Results
Communication Social Courtesies 0.75 0.73 0.58 0.67
@@
Problem-solving Test Review 0.89 0.92 0.79 0.70
@@ E.2 Fine-tuned Models Classification Results
Communication Social Courtesies 0.73 0.67 0.68
@@
Problem-solving Test Review 0.83 0.79 0.80
@@ Appendix F LIWC Results
Figures 4 and 5 present the co-occurrence of each strategy with the selected LIWC features.
@@
Figure 5: Co-Occurrence of Communication Strategies with Selected LIWC Features

### v2

@@ 6 Conclusion
This study highlights the potential of LLMs like GPT-4o for effective annotation of therapeutic dialogues. GPT-4o achieved a notable F1 score of 0.76, with a low entropy value of 0.035, demonstrating its efficiency and reliability in categorizing PST strategies. We fine-tuned transformer-based models on a sample of 5,000 therapist utterances annotated by GPT-4o and showed that ModernBERT exceeded the performance of our second-best LLM, GPT-4, showcasing the effectiveness of using LLM-generated annotations to fine-tune smaller, domain-specific models. We also expanded PST codebook to include a new Facilitators dimension, enriching our analysis of therapist-client interactions. We also analyzed Therapeutic Dynamics ( autonomy, self-disclosure, and metaphor) and linguistic patterns in our data to better understand ‘how’ PST strategies are used and in what contexts. Our research offers valuable insights into enhancing technology integration in mental health care, aiming to make therapeutic interventions both more accessible and more effective.
@@ 7 Limitations
Our study has several limitations, including the dependence on ‘transcribed’ therapy sessions, which may not fully capture the dynamic aspects of live interactions such as tone and non-verbal cues. The use of LLMs like GPT-4o and GPT-4 introduces potential biases from their internet-based training data, impacting the interpretation of therapeutic strategies. Furthermore, our analysis is solely based on English-language data, which limits the model’s applicability in non-English-speaking contexts. These limitations highlight the technology’s current stage and its focus on textual analysis, excluding the broader emotional and psychological facets of therapy sessions.
@@
Finally, we did not compare our results with any LLM-generated therapy data, which could offer a valuable comparison between real-world therapeutic interactions and simulated data produced by LLMs. Our focus is exclusively on the characteristics of authentic, real-world data from therapy sessions. Future research could explore this comparison, examining how well LLM-generated interactions align with or diverge from real-world therapeutic PST dialogues, potentially offering insights into improving the practicality and effectiveness of AI-driven interventions in mental health care.
@@ 8 Ethical Considerations
We are collaborating with a team of PST experts who were in charge of collecting and transcribing the data. We only have access to secondary data, and the primary data was collected as part of a larger task approved by the primary institution’s review board. As a result, we do not have access to any identifiable information about the clients or therapists. We understand that using anonymized therapy data raises concerns regarding privacy and the potential for re-identification of people involved in these sessions. We ensured that all data used in our analysis was thoroughly cleaned, with the help of our collaborators, and that strict measures were put in place to remove any remaining information from the transcriptions.
The application of LLMs in healthcare, particularly in sensitive areas like therapy, necessitates careful consideration of their implications. Ensuring responsible development and use of AI technologies in mental health care is crucial, requiring vigilance for biases, stakeholder engagement, and strong safeguards to maintain patient confidentiality and the integrity of the therapeutic process. This paper does not endorse the use of LLMs in therapeutic settings, nor does it claim that they are ready for such applications. Rather, our goal is to systematically evaluate the behavior of current LLMs when applied in therapeutic contexts. The use of AI in mental health care raises ethical concerns about automating therapy. AI tools, like LLMs, should assist rather than replace therapists, ensuring they do not undermine the personal, empathetic nature of therapy, which is crucial for effective treatment. While AI can provide useful insights, therapists must remain the primary decision-makers, as AI cannot fully grasp the complexities of human behavior and emotions as shown in our results and other related work. Accountability should always rest with human professionals to preserve the integrity and effectiveness of mental health care.
@@ Appendix B Therapists Strategies
Problem-solving Positive Mindset (Step One) Strategies building a positive approach towards challenges, emphasizing optimism and emotional awareness. “uh, they suggest trying to think of problems as challenges rather than threats.”
Defining Problems and Goals (Step Two) Setting achievable goals and identifying factual problems while recognizing obstacles. “Are there any obstacles to those goals? Can we list anything that is getting in the way?”
PS Core Generating Alternative Solutions (Step Three) Brainstorming solutions, differentiating between strategies, and encouraging open idea exploration. “So let’s brainstorm some specific strategies and tactics for how you can make that happen.”
Outcome Prediction and Planning (Step Four) Narrowing down solutions, assessing impacts, and creating an action plan. “Working through this worksheet…short-term personal consequences of self-care are pretty positive, right?”
Trying Out Solution Plan (Step Five) Implementing and assessing the solution strategy, troubleshooting, and acknowledging successes. “Then just monitoring your outcomes. Has the problem improved in some ways?”
@@
Facilitators Session Management* The therapist provides guidance on session logistics, manages time, keeps discussions on track, and helps plan future sessions. “So you got Skype up and running, and that is so fun!”
@@
Test Review* The therapist explains the PS Test and reviews the client’s results across five dimensions: positive/negative problem orientation, rational, impulsive, and avoidant problem-solving styles. “So positive and rational were your highest scores. Your positive was, uh, really, is almost as high as it could possibly get.”
@@ Appendix C Model Setting
Adam optimizer, learning-rate=2e-5, DeBERTa = microsoft/deberta-v3-base, MentalBERT = mental/mental-bert-base-uncased, FlanT5 = google/flan-t5-base ModernBERT = answerdotai/ModernBERT-base
@@ Appendix D Prompts
D.1 Classification of PST Strategies
D.1.1 Without Context
@@
Your Task: Select one strategy from the ‘PS core’ category and one strategy from the Facilitators category that best aligns with the therapist’s utterance. If none of the strategies are applicable, return ’None’ for that category.
@@
Facilitative Strategies: 6- Social Courtesies: Engaging in polite, yet somewhat superficial social interactions like greetings, small talk (e.g., asking about the weather or weekend plans), thanking the client for their time, and saying goodbye. 7- Session Management: Managing session logistics, such as informing the client about session length, guiding the direction of the conversation, refocusing when off-topic, or scheduling future sessions. 8- Therapeutic Engagement: Actively listening, validating the client’s feelings, providing emotional support, recalling previous information shared by the client, reflecting emotions, normalizing experiences, commenting on client strengths, offering personal disclosure, and offering comfort. 9- Test Review: Discussing the client’s Problem-Solving Test results, explaining the test, and reviewing five problem-solving dimensions: positive problem orientation, negative problem orientation, rational style, impulsive style, and avoidant style.
D.1.2 With Context
@@
Your Task: Select one strategy from the PS Core category and one strategy from the Facilitators category that best aligns with the therapist’s most recent utterance. You are also provided with the last two utterances: the therapist’s previous utterance and the client’s previous utterance. You should use only these two previous utterances as context, but for annotation, you should focus only on the most recent therapist’s utterance. If none of the strategies are applicable, return ’None’ for that category.
@@
D.2 Extracting Domain Specific Features
Prompt.
You are a helpful assistant tasked with analyzing a therapist’s utterances in a problem-solving therapy session. Your goal is to assess the utterances for the following aspects: 1. Autonomy-Supportive Language Output Labels: "Directive", "Non-Directive", or "N/A". Definitions: - Directive: - Focuses on the therapist’s goals, plans, or strategies for the client. - Therapist explicitly tells the client what to do or prescribes a solution. - Prioritizes achieving a specific outcome, often determined by the therapist. - May reflect the therapist’s beliefs, goals, or values rather than the client’s. - Frequently uses closed-ended or leading questions to guide the client toward a particular answer or decision. - Provides answers, advice, or logical consequences rather than encouraging exploration. - Non-Directive: - Focuses on the client’s goals, values, and intrinsic motivations. - Encourages the client to explore their thoughts and emotions freely without being led toward a specific conclusion. - Gives the client control over the direction of therapy and supports their decision-making process. - Avoiding judgment or imposition of therapist’s opinions. - Uses open-ended questions to encourage self-reflection and insight. - Attends to the client’s feelings and underlying experiences, fostering understanding and autonomy. - Accepts and explores resistance as part of the process, fostering collaboration to address challenges.
2. Self-Disclosure Output Labels: "Immediate", "Nonimmediate", or "N/A". Definitions: Self-disclosure refers to instances when the therapist intentionaly shares personal information, experiences, or feelings with the client. - Immediate Self-Disclosure: Sharing personal thoughts, feelings, or reactions related to the ongoing therapeutic interaction. - Nonimmediate Self-Disclosure: Discussing past experiences, personal anecdotes, or biographical information unrelated to the immediate session.
3. Open vs. Closed Questions Output Labels: "Open-Ended", "Closed-Ended", or "N/A". Definitions: - Open-Ended Questions: Questions that encourage detailed, thoughtful, and expansive responses. These questions often begin with "how," "what," "why," or "tell me about" and invite the client to explore their thoughts and feelings freely. - Closed-Ended Questions: Questions that elicit specific, often short responses such as "yes," "no," or concise information. These questions are structured and limit the scope of the client’s answer. N/A: Use this label if the utterance does not fit into either category.
4. Metaphors: Output Labels: "yes, <reasoning for metaphor detection>" (for use of metaphor) or "no" (for no use of metaphor). s Definition: According to conceptual metaphor theory, metaphor facilitates a mapping of attributes or characteristics from source domain to target domain. For example, the word “invested” in the sentence “I have invested a lot of time in her” is a metaphorical expression. The source domain implied by this metaphor is the domain of money and the target domain implied by this metaphor is the domain of time. Based on the above information, decide whether this utterance uses metaphor in it and mention the phrase, source and domain of the metaphor.
Required Output Format (JSON):
"autonomy": "<string>", "self disclosure": <integer>, "question type": "<string>", "metaphor": "<string>"
@@ E.1 LLM Classification Results
Facilitators Social Courtesies 0.75 0.73 0.58 0.67
@@
Test Review 0.89 0.92 0.79 0.70
@@ E.2 Fine-tuned Models Classification Results
Facilitators Social Courtesies 0.73 0.67 0.68
@@
Test Review 0.83 0.79 0.80
@@ Appendix F LIWC Results
Figures 8 and 9 present the co-occurrence of each strategy with the selected LIWC features.
@@
Figure 9: Co-Occurrence of Facilitative Strategies with Selected LIWC Features
Appendix G Question Type vs. Autonomy
Figure 10 shows the co-occurrence matrix between question type and autonomy labels.
Figure 10: Co-occurrance of Question Type and Autonomy Labels